brandonHollinger:

Solve for x: |4x + 12| = 16 how would you do your work for this?

mxddi3:

you need to get the 4x by itself so you subtract 12 from each side, giving you 4x=4 then you just divide @brandonHollinger

brandonHollinger:

thank you

Shadow:

\[|4x + 12 | = 16\] The way you solve for these is by writing two equations: \[4x + 12 = 16\] \[4x + 12 = -16\] For the first, begin by subtracting 12 from both sides \[4x = 4\] Divide by 4 \[x = 1\] For the second equation, subtract 12 from both sides \[4x = -28\] Divide by 4 \[x = -7\]